# Infinite-Horizon Graph Filters: Leveraging Power Series to Enhance Sparse Information Aggregation

### v2

@@ 1 Introduction
i) Long-range Dependencies: Balancing the trade-off between the receptive field size and feature distinctiveness is a challenging aspect in GNNs. On the one hand, deeper GNNs Eli Chien and Milenkovic (2021); Li et al. (2019); He et al. (2021) offer a larger receptive field, which allows for the incorporation of information from a broader range of the graph. However, this comes with the downside of feature homogenization across nodes, leading to the phenomenon known as over-smoothing. On the other hand, GNNs with shallower structures Rusch et al. (2023), while avoiding over-smoothing, face limitations in capturing long-range dependencies due to their smaller receptive field. This limitation is particularly significant in real-world graphs, such as social networks or protein interaction networks, where understanding distant node relationships is crucial. ii) Graph Sparsity: A sparse graph is a type of graph in which the number of edges is relatively low compared to the total number of possible edges. Such graphs are common in real-world scenarios. For instance, Berger et al. (2005); Ahad N. et al. (2023) reveals that the degree distribution of real-world graphs typically follows a power-law distribution. In particular, the citation network-based Cora and Citeseer datasets exhibit remarkable levels of sparsity, measured at 99.93% and 99.96%, respectively. Because sparse graphs contain less explicit information due to fewer edges, it is hard to mine effective representations, even for contemporary GNNs. As shown in Figure 1, as the number of edges decreases gradually, the performance of GNNs drops correspondingly, indicating that learning effective representations on sparse graphs remains an unresolved challenge.
@@ 2 Related Work
Polynomial filters encompass ResGCN Li et al. (2019), BernNet He et al. (2021), and GPR-GNN Eli Chien and Milenkovic (2021), etc.. ResGCN employs residual connections to facilitate information transfer between different layers, alleviating the over-smoothing issues. However, ResGCN focuses more on intermediate information, neglecting the significance of proximal information. BernNet utilizes Bernstein polynomials to aggregate information across different layers. Nevertheless, due to layer constraints, BernNet struggles to extend its reach to more distant information, and parameter learning becomes more challenging. APPNP Gasteiger et al. (2022) and GRAND Feng et al. (2022) utilize feature propagation but within a limited number of hops. GPR-GNN unifies the representation of parameters between different layers in a general polynomial formula, therefore APPNP, ResGCN and BernNet can all be regarded as special case of GPR-GNN. However, GPR-GNN requires learning relative control parameters between layers. Similar to BernNet, the receptive field of view is limited, and parameter learning poses a significant challenge.
@@ 4.1.2 Power Series Graph Filters
Katz comb-pass $\frac{1}{\beta_{a}}[(I-\beta_{a}\hat{A})^{-1}-I]$ ${\beta_{a}}^{n-1}$ ${\tilde{A}_{sym}}$ +$\infty$+ ∞
@@ 5.1.1 Datasets
  • Citeseer3: Similar to the Cora dataset, Citeseer is a citation graph-based node classification dataset.
@@ 5.1.2 Baselines
[…]
@@ 6 Conclusion
This paper focuses on the design of power series-enhanced GNNs to address the challenges of long-range dependencies and the sparse graph. To ensure the efficiency of our GPFN, a graph filter using convergent power series from the spectral domain is introduced in this paper. The effectiveness of our GPFN is verified by theoretical analysis from both spectral and spatial perspectives and experimental results, demonstrating its superiority over state-of-the-art graph learning techniques on benchmark datasets. Future directions for investigation include exploring diverse filters such as the mid-pass filter and integrating the diffusion model to further bring the explanation into our GPFN.
@@ Appendix C Baselines
[…]

### v3

@@ 1 Introduction
i) Long-range Dependencies: Balancing the trade-off between the receptive field size and feature distinctiveness is a challenging aspect in GNNs. On the one hand, deeper GNNs Eli Chien and Milenkovic (2021); Li et al. (2019); He et al. (2021) offer a larger receptive field, which allows for the incorporation of information from a broader range of the graph. However, this comes with the downside of feature homogenization across nodes, leading to the phenomenon known as over-smoothing. On the other hand, GNNs with shallower structures Rusch et al. (2023), while avoiding over-smoothing, face limitations in capturing long-range dependencies due to their smaller receptive field. This limitation is particularly significant in real-world graphs, such as social networks or protein interaction networks, where understanding distant node relationships is crucial.
ii) Graph Sparsity: A sparse graph is a type of graph in which the number of edges is relatively low compared to the total number of possible edges. Such graphs are common in real-world scenarios. For instance, Berger et al. (2005); Ahad N. et al. (2023) reveals that the degree distribution of real-world graphs typically follows a power-law distribution. In particular, the citation network-based Cora and Citeseer datasets exhibit remarkable levels of sparsity, measured at 99.93% and 99.96%, respectively. Because sparse graphs contain less explicit information due to fewer edges, it is hard to mine effective representations, even for contemporary GNNs. As shown in Figure 1, as the number of edges decreases gradually, the performance of GNNs drops correspondingly, indicating that learning effective representations on sparse graphs remains an unresolved challenge.
@@ 2 Related Work
Polynomial filters encompass ResGCN Li et al. (2019) (including those GNNs employing long-range residual connection such as Graph Transformer Wu et al. (2021)), BernNet He et al. (2021), and GPR-GNN Eli Chien and Milenkovic (2021), etc.. ResGCN employs residual connections to facilitate information transfer between different layers, alleviating the over-smoothing issues. However, ResGCN focuses more on intermediate information, neglecting the significance of proximal information. BernNet utilizes Bernstein polynomials to aggregate information across different layers. Nevertheless, due to layer constraints, BernNet struggles to extend its reach to more distant information, and parameter learning becomes more challenging. APPNP Gasteiger et al. (2022) and GRAND Feng et al. (2022) utilize feature propagation but within a limited number of hops. GPR-GNN unifies the representation of parameters between different layers in a general polynomial formula, therefore APPNP, ResGCN and BernNet can all be regarded as special case of GPR-GNN. However, GPR-GNN requires learning relative control parameters between layers. Similar to BernNet, the receptive field of view is limited, and parameter learning poses a significant challenge.
@@ 4.1.2 Power Series Graph Filters
Katz comb-pass $[(I-\beta_{a}\hat{A})^{-1}-I]/{\beta_{a}}$ ${\beta_{a}}^{n-1}$ ${\tilde{A}_{sym}}$ +$\infty$+ ∞
@@ 5.1.1 Datasets
  • Citeseer5: Similar to the Cora dataset, Citeseer is a citation graph-based node classification dataset.
@@ 5.1.2 Baselines
We compare our GPFN  with $18$ baselines, from four MR categories for comprehensive experiments: i) Non-graph filter-based methods: MLP Rosenblatt (1963) and LP Zhu and Ghahramani (2002). ii) Monomial graph filter-based methods: GCN Kipf and Welling (2016), GAT Veličković et al. (2018), GIN Xu et al. (2019), AGE Cui et al. (2020), GCN-SGC and GAT-SGC Wu et al. (2019). iii) Polynomial graph filter-based methods: ChebGCN Defferrard et al. (2016), GPR-GNN Eli Chien and Milenkovic (2021), APPNP Gasteiger et al. (2022), BernNet He et al. (2021), GRAND Feng et al. (2022), GCNII Chen et al. (2020), ADC Zhao et al. (2021), DGC Wang et al. (2021), D2PT Liu et al. (2023), HiGNN Huang et al. (2023), HiD-GCN Li et al. (2024) Res-GCN and Res-GAT Li et al. (2019). Detailed description of baselines can be referred to in Appendix C.
@@ 6 Conclusion
This paper focuses on the design of power series-enhanced GNNs to address the challenges of long-range dependencies and sparse graphs. To ensure the efficiency of our GPFN, a graph filter using convergent power series from the spectral domain is introduced in this paper. The effectiveness of our GPFN is verified by theoretical analysis from both spectral and spatial perspectives and experimental results, demonstrating its superiority over state-of-the-art graph learning techniques on benchmark datasets. Future directions for investigation include exploring diverse filters such as the mid-pass filter and integrating the diffusion model to further bring the explanation into our GPFN. Moreover, we would like to deeper explore and analysis GPFN on heterogeneous graphs.
@@ Appendix C Baselines
GCNII https://github.com/chennnM/GCNII
ADC https://github.com/abcbdf/ADC
DGC https://github.com/yifeiwang77/DGC
@@
  • GCNII Chen et al. [2020]: GCNII is an extension of the vanilla GCN model with two simple yet effective techniques– Initial residual and Identity mapping.
  • ADC Zhao et al. [2021]: ADC learns a dedicated propagation neighborhood for each GNN layer and each feature channel, making the GNN architecture fully coupled with graph structures—the unique property that differs GNNs from traditional neural networks.
  • DGC Wang et al. [2021]: DGC decouples the terminal time and the feature propagation steps, making it more flexible and capable of exploiting a very large number of feature propagation steps.
@@
  • D2PT Liu et al. [2023]: D2PT performs the dual-channel diffusion message passing with the contrastive-enhanced global graph information on the sparse graph.
@@ Appendix D Hyper-parameter Settings of Baselines
Appendix E Training Time Comparison
Table 5: Training time comparison between GPFN and baselines on Cora, Citeseer and AmaComp datasets with masking ratios (MR) equals 0.
Avg. Training Time (ms) / Epoch
Dataset  Cora  Citeseer  AmaComp
Baselines GCN  25.64  25.93  24.41
GCN-SGC  25.79  29.33  48.57
APPNP  35.14  32.40  41.69
GRAND  32.18  34.89  53.04
GPRGNN  30.15  27.75  24.86
BernNet  25.79  29.33  48.57
HiGCN  56.89  59.44  72.13
HiD-GCN  76.53  86.22  92.09
GCNII  91.62  91.63  86.95
GPFN GCN-Katz  27.71  29.75  27.44
GCN-S2  29.89  26.64  32.52
GCN-S3  30.03  26.87  35.92
We compare our GPFN with some SOTA methods that address over-smoothing problems, and the runtime per epoch is shown in Table 5. It is worth noting that despite our methods making the graph denser, the computational time remains relatively unchanged compared to other methods. GPFN keeps a better balance between time and performance.
Appendix F Complexity Analysis
Due to matrix eigenvalue decomposition and matrix inverse operations involved in our computation, the time complexity of the entire process is $O(N^{3})$, this is the same with GPRGNN, HiGCN, HiD-GCN and GCNII. However, our GPFN takes into account long-range dependencies, achieving excellent performance without the need for deeper GNN layers, significantly reducing the model’s parameter. In contrast, GCNII requires 64 layers to achieve the best performance on Cora.
Appendix G Motivation Explanation
The main motivation of our paper lies in addressing two critical challenges faced by existing GNNs: handling long-range dependencies and mitigating the adverse impacts of graph sparsity. Motivated by the power function, we provide a solution that effectively captures long-range dependencies while leveraging the sparse nature of real-world graphs to enhance representation learning. The proposed GPFN framework is analyzed from both spectral and spatial domains. In the spectral domain, it serves as a flexible graph filter framework capable of accommodating different filter types. In the spatial domain, it acts as an infinite information aggregator, leveraging power series filters to aggregate neighborhood information across an infinite number of hops. This enables the enlargement of the graph’s receptive field, thereby addressing the challenge of capturing long-range dependencies effectively.